# LLM-Personalize: Aligning LLM Planners with Human Preferences via Reinforced Self-Training for Housekeeping Robots

### v1

@@ III-D1 Imitation Learning
$\mathcal{L}_{NLL}=-\mathbb{E}_{(\mathbf{x},\mathbf{y})\in\mathcal{D_{\textnormal{demo}}}}\left[\sum_{\tau=1}^{|\mathbf{y}|}\log P_{\theta}(\mathbf{y}_{\tau}\mid\mathbf{y}_{1:\tau-1},\mathbf{x})\right]$ (1)
The above demonstration design guides the LLM planner to more accurately extract information from complex input contexts and improve plan executability. More importantly, it ensures that each plan has a uniform objective (i.e, perform exploration or rearrangement of a single object), allowing the plan to be straightforwardly annotated with user preferences.
III-D2 Iterative Self-training
Next, we fine-tune the bootstrapped LLM planner to further improve personalization via iterative self-training on a set of training tasks. This allows the LLM planner to explore more rearrangement options and improve its placement decisions through imitating the positive examples. As shown in Fig. 3, to start a self-training iteration, we use the LLM planner to explore (with randomness by adjusting the temperature parameter) by collecting episodes of experiences on the training tasks and we log $M$ interactions as tuples of prompt, response and outcomes: $\{(\mathbf{x}^{i},\mathbf{y}^{i},out^{i})\}_{i=1}^{M}$ (Example outcome: pan 1 moved from kitchen 0 table 6 to kitchen 0 counter 1). We annotate each prompt and response pair $(\mathbf{x}^{i},\mathbf{y}^{i})$ with the reward $r^{i}\in\{1,0,-1\}$ for the rearrangement outcomes according to the user preferences. Then, we collect a self-training dataset by picking the prompt and response pairs with positive rewards $\mathcal{D}_{\textnormal{self-train}}=\{(\mathbf{x}^{i},\mathbf{y}^{i})\mid r^{i}>0\}_{i=1}^{M}$ The final step of this self-training iteration is to perform supervised fine-tuning of the bootstrapped LLM model over $\mathcal{D}_{\textnormal{self-train}}$ with the NLL objective as defined in Equation (1). This procedure is repeated iteratively where each self-training iteration performs interaction collection and fine-tuning over the LLM obtained from the previous iteration.
IV EXPERIMENTS
@@
TABLE I: Average success rate on train and test sets across scenes. Each entry denotes the mean ± standard error of the mean across episodes. (Boldface: best variant across the task set, ST: self-training, IL: imitation learning)
Scene ID Scene 1 Scene 2 Scene 3 Scene 4 Average
Task Set train test train test train test train test test
Baselines SayCan -2.6 ± 1.9 0.0 ± 0.0 -1.2 ± 1.2 -10.6 ± 6.8 -3.3 ± 2.2 -8.0 ± 4.9 -1.6 ± 1.6 0.0 ± 0.0 -4.6 ± 2.9
SayPlan -7.0 ± 5.9 -5.0 ± 5.0 -6.8 ± 2.7 -1.6 ± 9.2 0.4 ± 4.5 -13.0 ± 8.3 -10.7 ± 3.8 -12.3 ± 15.3 -7.9 ± 9.4
LLM-Planner 5.3 ± 4.4 -3.6 ± 4.8 -8.4 ± 3.8 -9.8 ± 6.3 -14.2 ± 4.0 -4.0 ± 3.3 -29.6 ± 5.2 -30.2 ± 4.8 -11.8 ± 4.8
Ours LLM-Personalize (IL) 4.1 ± 2.6 17.6 ± 6.1 -3.3 ± 3.0 12.6 ± 9.0 22.6 ± 2.6 24.3 ± 4.4 10.9 ± 3.3 25.7 ± 6.7 20.0 ± 6.5
LLM-Personalize (ST iter=1) 17.9 ± 3.7 25.8 ± 6.6 19.4 ± 2.9 21.7 ± 5.6 32.4 ± 3.2 41.4 ± 6.3 24.2 ± 3.2 10.2 ± 6.7 24.7 ± 6.3
LLM-Personalize (ST iter=2) 25.5 ± 3.1 29.6 ± 5.4 18.5 ± 2.6 25.2 ± 4.0 33.5 ± 3.8 43.3 ± 4.3 29.1 ± 2.9 20.4 ± 6.8 29.6 ± 5.1
IV-A Overview
We empirically evaluate the performance of LLM-Personalize on the Housekeep benchmark. In particular, we aim to evaluate the hypothesis: Optimizing LLM planner through imitation learning and self-training allows the LLM planner to improve planning performance and alignment with user preferences. We demonstrate this hypothesis through improved rearrangement success rate compared with baseline LLM planners and provide qualitative results that showcase the plans generated. Our ablations studies further evaluate plan executability, exploration and cross-domain (scene) generalisation of LLM-Personalize in the different training phases. These analyses provide a deeper dive into the improved task success and offer more insights for personalizing LLM planner in future applications.
IV-B Experiment Setup
For our main results, we evaluate LLM-Personalize on in-domain (scene) adaptation performance on 4 different scenes in the Housekeep benchmark.
IV-B1 Environment setup
Each scene contains a different layout of rooms and receptacles, as shown in Fig. 4. A task is instantiated with a random selection of $5-10$ objects placed on different receptacles, among them $\sim 3-7$ objects are misplaced and needs to be rearranged. For each object, there is a list of correct receptacles not known to the agent. The task is challenging since the agent needs to explore the house, identify misplaced objects and their correct receptacles, and avoid moving correctly placed objects to wrong receptacles.
IV-B2 Evaluation metrics
We evaluate the improvement of success rate of in-domain adaptation. The success rate [10] is defined as the percentage of misplaced objects that are correctly re-arranged at the end of the task, among all misplaced objects at the start of the task.
Success Rate=(#correct at the end#correct at the start#total misplaced objects at the $\textit{\scriptsize Success Rate}=\left(\frac{\#\textit{correct at the end}-\#\textit{correct at the start}}{\#\textit{total misplaced objects at the start}}\right)$
@@ IV-B3 Architecture and Baselines
We compare LLM-Personalize with state-of-the-art LLM planning baseline methods: LLM-Planner [1], SayPlan [3] and SayCan [2]. We use GPT-3.5-turbo as the LLM across all compared methods, and the temperature during LLM generation is set to $1$ to reduce deterministic repetition in LLM responses (default range is $0-2$). For all methods, the prompt includes the instruction, graph description and two examples. We allow all methods to plan iteratively after the robot executed the previous plan. LLM-Planner with these configurations is adopted as the base model of LLM-Personalize, which we then optimize using imitation learning and self-training. The fine-tuning as defined in Eq.(1) is performed via the OpenAI fine-tune API. To adapt SayCan to Housekeep with large number of available actions, we adopt the implementation in [1], where the list of affordable high-level actions (e.g., go to a discovered object) are provided in each prompt, together with the list of previously executed actions. For SayPlan, we additionally provide the state and affordance of the receptacles and objects (e.g., pick up) as in [3] and allow 10 LLM semantic search steps and 5 revision (re-plan) retries for each plan iteration where the final revised plan is executed, and the revision feedback is provided by a verification module (e.g., cannot pick up obj as the agent is holding another object).
@@
Figure 5: Demonstration of four planning iterations generated and executed by LLM-Personalize (top row) and the resulting graphs (bottom row) on a test task in Housekeep. Green/red object (leaf) nodes indicate correct/wrong placements. The object being moved is shown in boldface with highlighted edge. This episode starts with 2 correctly placed objects and 5 misplaced objects (left), and changed to 6 correctly placed objects and only 1 misplaced objects after rearrangements (right). For clarity, the graphs only show receptacles with objects and omit all other receptacles.
IV-C Main Results
IV-C1 Quantitative Results
@@
We compare the performance in terms of average success rate of LLM-Personalize with the baseline methods in Table I, and we also demonstrate the effectiveness of our optimization framework by comparing the performance of LLM-Personalize at different optimization phases, namely, the base version (using LLM-Planner), imitation learning (IL), and various iterations of self-training (ST). Each table entry shows the mean and standard error of the mean across all tasks in the train/test set and 5 runs per task. Overall, we observe that LLM-Personalize significantly outperforms all baseline methods in terms of mean success rate across the tested scenes. For example, on the test set of Scene 1, the success rate of the baselines are near zero or negative, while LLM-Personalize achieved $29.6\%$ after imitation learning and two self-training iterations. This trend can be similarly observed across all scenes. To examine closely into the difference, we analysed some detailed prompt and responses and identified that SayCan often has difficulty picking the best (high-level) action from a large number of available actions due to the large number of objects, receptacles and rooms. LLM-Planner is often able to produce correct pick and place action sequences following the in-context examples. Compared with LLM-Planner, SayPlan improves slightly (on 5 out of 8 task sets) as a result of improvement in plan executability due to revision with feedback. However, all baselines have difficulty knowing whether an object is misplaced or correctly placed, as well as the correct receptacles to place objects, due to lack of personalization. As a result, the negative scores across the baseline methods are often due to picking and placing correctly placed objects onto wrong receptacles.
@@
TABLE II: Ablation Study: Cross-domain generalisation success rate. (Boldface: best across the task set, ST1/2: self-training iteration 1 or 2, IL: imitation learning)
@@
Comparing the different stages of LLM-Personalize on the test sets, we observe a general trend where the combination of imitation learning and self-training lead to better results. First, bootstrapping from demonstrations improves over LLM-Planner (i.e., the base LLM-Personalize model). For example, on Scene 1, the success rate improved from $-3.6\%$ to $17.6\%$. This improvement is a result of improved executability due to action sequencing, better context understanding (e.g., agent correctly extracts and uses object names from the prompt), and initial alignment to personalized preferences shown in the demonstrations. Second, the self-training iterations further improves from the bootstrap variants with improved alignment with personalized preferences. For example, on Scene 1, the success rate after two iteration of self-training grows from $17.6\%$ to $29.6\%$ compared with the imitation bootstrapped variant. We also observed that after bootstrapping, the model learns to explore object placements with improved accuracy during self-training and with more self-training iterations, the model exploits and commits to the learned correct placements while avoiding wrong ones.
Comparing the train and test performances of LLM-Personalize for each scene, we observe that the testing performance generally increases with increased training performance, except for Scene 3, where the test performance drops as a result of overfitting. This shows that our model is able to learn personalized preferences seen during training, and generalize to unseen object combinations and placements.
IV-C2 Qualitative Results
In Fig. 5, we present plans generated by LLM-Personalize and executed by the robot on a test task in the Housekeep simulator and the resulting graph of the household scene after each plan iteration. We can see that the agent learned to start by exploring the house, then rearrange one misplaced object at a plan iteration, and successfully rearranged 4 out of 5 misplaced objects.
IV-D Ablation Studies
(a) Executability
(b) Unique Placements
Figure 6: Ablation study: (a) shows the percentage of executable high-level action steps, (b) shows the unique placements executed. x-axis refers to LLM-Personalize at different optimization phases – Base: before optimization, IL: imitation learning, ST: self-training. Each point is an average value over 25 episodes (5 runs per task over 5 tasks in the test set) and shaded area refers to standard error of the mean.
IV-D1 Cross-domain Transfer Results
In addition to the improved in-domain adaptation results in Table I, we show in Table II cross-domain transfer performance of LLM-Personalize. In this experiment, we train the model on a source scene (e.g., scene 2 or scene 4), and observe the performance change on the test set on a different scene (e.g., scene 1 or scene 3) with different rooms and receptacles. From the table, we can observe that through imitation learning and self-training, LLM-Personalize is able to transfer to a different scene with improved test performance.
IV-D2 Executability
In Fig. 6(a) we present the executability improvement of LLM-Personalize (IL, ST1, ST2) compared to the base LLM-Planner (Base). Each point refers to the average percentage of high-level actions generated by the LLM planner that are successfully executed per episode. As expected the IL bootstrapping significantly improved the planner’s executability, due to improved context (prompt) understanding and action sequencing, enabling LLM-Personalize to produce high quality training examples for self-training, and we can observe consistently high executability from the LLM-Personalize(ST1) and LLM-Personalize(ST2) variants.
IV-D3 Exploration vs. Exploitation
To analyse how the agent’s exploration vs. exploitation behavior changes at different optimization phases, we show in Fig. 6(b) the average number of unique placements executed per episode, where each unique placement refers to a pair of object and receptacle where the agent placed the object on the receptacle. Higher degree of exploration behavior is indicated by higher number of unique placements, and in contrary, lower number of unique placements indicates more exploitation behavior. As shown in Fig. 6(b), we see an increasing trend in exploration from LLM-Planner (Base) to LLM-Personalize(IL), partly due to improved plan executability. From IL to ST1, the agent further increased exploration, where it explores placements beyond the behaviors learned from demonstrations. For example, on Scene 3 the average unique placements increased from 2.32 to 9.48. Compared with ST1, ST2 typically shows higher exploitation behavior where the agent learns to commit to the correct placement combinations for better task performance.
V CONCLUSIONS
@@
We proposed LLM-Personalize, a household robotics agent framework with an LLM-based architecture capable of performing long-horizon planning in multi-room, partially observable household scenarios, and an optimisation pipeline designed to personalize the LLM planner according to user preferences. This novel approach effectively addresses the critical gap in the personalization of LLM planners for household robotics. Our model achieves superior alignment with user preferences, outperforming existing work in the challenging Housekeep rearrangement tasks. Looking forward, the versatility and scalability of the agent design and optimization pipeline makes LLM-Personalize a promising solution for a broad range of household robotics applications.

### v2

@@ 3.3.1 Imitation Learning
$\mathcal{L}_{NLL}=-\mathbb{E}_{(\mathbf{x},\mathbf{y})\in\mathcal{D_{\textnormal{demo}}}}\left[\sum_{\tau=1}^{|\mathbf{y}|}\log P_{\theta}(\mathbf{y}_{\tau}\mid\mathbf{y}_{1:\tau-1},\mathbf{x})\right]$
The above demonstration design guides the LLM planner to more accurately extract information from complex input contexts and improve plan executability. More importantly, it ensures that each plan has a uniform objective (i.e, perform exploration or rearrange a single object), allowing the plan to be straightforwardly annotated with user preferences during Self-Training.
3.3.2 Iterative Reinforced Self-Training
Next, we fine-tune the bootstrapped LLM planner to further improve personalization via iterative Self-Training on a set of training tasks. This allows the LLM planner to explore more rearrangement options and improve its placement decisions through imitating the positive examples. As shown in Fig. 3, to start a Self-Training iteration, we use the LLM planner to explore by collecting episodes of experiences on the training tasks and we log $M$ interactions as tuples of prompt, response and outcomes: $\{(\mathbf{x}^{i},\mathbf{y}^{i},out^{i})\}_{i=1}^{M}$ (Example outcome: pan 1 moved from kitchen 0 table 6 to kitchen 0 counter 1). We annotate each prompt and response pair $(\mathbf{x}^{i},\mathbf{y}^{i})$ with the reward $r^{i}\in\{1,0,-1\}$ for the rearrangement outcomes according to the user preferences. Then, we collect a Self-Training dataset by picking the prompt and response pairs with positive rewards $\mathcal{D}_{\textnormal{self-train}}=\{(\mathbf{x}^{i},\mathbf{y}^{i})\mid r^{i}>0\}_{i=1}^{M}$ The final step of this Self-Training iteration is to perform supervised fine-tuning of the bootstrapped LLM model over $\mathcal{D}_{\textnormal{self-train}}$ with the NLL objective as defined in Equation (1). This procedure is repeated iteratively where each Self-Training iteration performs interaction collection and fine-tuning over the LLM obtained from the previous iteration.
4 EXPERIMENTS
@@
Table 1: Average success rate on train and test sets across scenes. Each entry denotes the mean ± standard error of the mean across episodes. (Boldface: best variant across the task set, ST: self-training, IL: imitation learning)
Scene ID Scene 1 Scene 2 Scene 3 Scene 4
Task Set train test train test train test train test
Baselines SayCan -2.6 ± 1.9 0.0 ± 0.0 -1.2 ± 1.2 -10.6 ± 6.8 -3.3 ± 2.2 -8.0 ± 4.9 -1.6 ± 1.6 0.0 ± 0.0
SayPlan -7.0 ± 5.9 -5.0 ± 5.0 -6.8 ± 2.7 -1.6 ± 9.2 0.4 ± 4.5 -13.0 ± 8.3 -10.7 ± 3.8 -12.3 ± 15.3
LLM-Planner 5.3 ± 4.4 -3.6 ± 4.8 -8.4 ± 3.8 -9.8 ± 6.3 -14.2 ± 4.0 -4.0 ± 3.3 -29.6 ± 5.2 -30.2 ± 4.8
Ours LLM-Personalize (IL) 4.1 ± 2.6 17.6 ± 6.1 -3.3 ± 3.0 12.6 ± 9.0 22.6 ± 2.6 24.3 ± 4.4 10.9 ± 3.3 25.7 ± 6.7
LLM-Personalize (ST iter=1) 17.9 ± 3.7 25.8 ± 6.6 19.4 ± 2.9 21.7 ± 5.6 32.4 ± 3.2 41.4 ± 6.3 24.2 ± 3.2 10.2 ± 6.7
LLM-Personalize (ST iter=2) 25.5 ± 3.1 29.6 ± 5.4 18.5 ± 2.6 25.2 ± 4.0 33.5 ± 3.8 43.3 ± 4.3 29.1 ± 2.9 20.4 ± 6.8
We aim to evaluate the hypothesis: Optimizing LLM planner through imitation learning and Self-Training allows the LLM planner to improve planning performance and alignment with user preferences. We demonstrate this through improved rearrangement success rate compared with baseline LLM planners and provide qualitative results that showcase the plans generated. Our ablations studies further evaluate plan executability, exploration and cross-domain (scene) generalisation of LLM-Personalize in the different training phases.
4.1 Experiment Setup
We evaluate LLM-Personalize on 4 different scenes in Housekeep, each featuring a unique layout of rooms and receptacles, as shown in Fig. 4. A task is instantiated with a random selection of $5-10$ objects placed on different receptacles, among them $\sim 3-7$ objects are misplaced and needs to be rearranged. For each object, there is a list of correct receptacles not known to the agent. The task is challenging as the agent must explore the house, identify misplaced objects and their correct receptacles, and avoid removing correctly placed objects.
Evaluation metrics: We evaluate the agent performance via success rate Kant et al. (2022), defined as the percentage of misplaced objects that are correctly re-arranged at the end of the task, among all misplaced objects at the start of the task.
Success Rate=(#correct at the end#correct at the start#total misplaced objects at the $\textit{Success Rate}=\left(\frac{\#\textit{correct at the end}-\#\textit{correct at the start}}{\#\textit{total misplaced objects at the start}}\right)$
@@ 4.1 Experiment Setup
Architecture and Baselines We compare LLM-Personalize with the baseline LLM planning methods: LLM-Planner Song et al. (2023), SayPlan Rana et al. (2023) and SayCan Ahn et al. (2022). For all methods, we use GPT-3.5-turbo with temperature set to $1$ to reduce deterministic repetition in LLM responses (default range is $0-2$). The prompt includes the instruction, graph description and two examples. We allow all methods to plan iteratively after the robot executed the previous plan. LLM-Planner with these configurations is adopted as the base model of LLM-Personalize, which we then optimize using imitation learning and Self-Training. The fine-tuning as defined in Eq.(1) is performed via the OpenAI fine-tune API. For more details please refer to Appendix A.3.
@@
Figure 5: Demonstration of four planning iterations generated and executed by LLM-Personalize (top row) and the resulting graphs (bottom row) on a test task in Housekeep. Green/red object (leaf) nodes indicate correct/wrong placements. The object being moved is shown in boldface. This episode starts with 2 correctly placed objects and 5 misplaced objects (left), and changed to 6 correctly placed objects and only 1 misplaced objects after rearrangements (right). For clarity, the graphs only show receptacles with objects and omit all other receptacles.
4.2 Main Results
@@
Quantitative Results We compare the average success rate of LLM-Personalize with the baselines in Table 1, and we show the effectiveness of our optimization framework by comparing the performance of LLM-Personalize at different optimization phases, namely, the base version (using LLM-Planner), imitation learning (IL), and various iterations of Self-Training (ST). Each table entry shows the mean and standard error of the mean across all tasks in the train/test set and 5 runs per task. Overall, we observe that LLM-Personalize significantly outperforms all baseline methods across the tested scenes. For example, on the test set of Scene 1, the success rate of the baselines are near zero or negative, while LLM-Personalize achieved $29.6\%$ after imitation learning and two Self-Training iterations. Similar trends can be observed across all scenes. After examining some detailed prompts and responses we identified that SayCan often has difficulty picking the best high-level action from a large number of available actions due to the large number of objects, receptacles and rooms. LLM-Planner is often able to produce correct pick and place action sequences following the in-context examples. Compared with LLM-Planner, SayPlan improves slightly (on 5 out of 8 task sets) as a result of improvement in plan executability due to revision with feedback. However, all baselines have difficulty knowing whether an object is misplaced or correctly placed, as well as the correct receptacles to place objects, due to lack of personalization. As a result, the negative scores across the baseline methods are often due to picking and placing correctly placed objects onto wrong receptacles.
Comparing the different stages of LLM-Personalize on the test sets, we observe a general trend where the combination of imitation learning and Self-Training lead to better results. First, bootstrapping from demonstrations improves over LLM-Planner (i.e., the base LLM-Personalize model). For example, on Scene 1, the success rate improved from $-3.6\%$ to $17.6\%$. This improvement is a result of improved executability due to action sequencing, better context understanding (e.g., agent correctly extracts and uses object names from the prompt), and initial alignment to personalized preferences shown in the demonstrations. Second, the Self-Training iterations further improves performance with improved alignment with personalized preferences. For example, on Scene 1, the success rate after two iteration of Self-Training grows from $17.6\%$ to $29.6\%$ compared with the imitation bootstrapped variant. We also observed that after bootstrapping, the model learns to explore object placements with improved accuracy during Self-Training and with more Self-Training iterations, the model exploits and commits to the learned correct placements while avoiding wrong ones.
Comparing the train and test performances of LLM-Personalize, we observe that the testing performance generally increases with increased training performance, except for Scene 3, where the test performance drops as a result of overfitting. This shows that our model is able to learn personalized preferences seen during training, and generalize to unseen object combinations and placements.
Qualitative Results In Fig. 5, we present plans generated by LLM-Personalize and executed by the robot on a test task in Housekeep and the resulting graph of the household scene after each plan iteration. We can see that the agent learned to start by exploring the house, then rearrange one misplaced object at a plan iteration, and successfully rearranged 4 out of 5 misplaced objects.
4.3 Ablation Studies
(a) Executability
(b) Unique Placements
Figure 6: Ablations (a) percentage of executable high-level actions (b) unique placements executed. x-axis refers to LLM-Personalize at different phases – Base: before optimization, IL: imitation learning, ST: Self-Training. Each point is an average value over 25 episodes (5 runs per task over 5 tasks in the test set) and shaded area refers to standard error of the mean.
Cross-domain Transfer Results In addition to the improved in-domain adaptation results in Table 1, we show in Table 2 cross-domain transfer performance of LLM-Personalize. In this experiment, we train the model on a source scene (e.g., scene 2), and observe the performance change on the test set on a different scene (e.g., scene 1) with different rooms and receptacles. From the table, we can observe that through imitation learning and Self-Training, LLM-Personalize is able to transfer to a different scene with improved test performance.
Executability In Fig. 6(a) we present the executability improvement of LLM-Personalize (IL, ST1, ST2) compared to the base LLM-Planner (Base). Each point refers to the average percentage of high-level actions generated by the LLM planner that are successfully executed per task. The IL bootstrapping significantly improved the planner’s executability, due to improved context understanding and action sequencing, enabling LLM-Personalize to produce high quality training examples for Self-Training, and we can observe consistently high executability from the LLM-Personalize(ST1) and LLM-Personalize(ST2) variants.
Exploration vs. Exploitation To analyse how the agent’s exploration vs. exploitation behavior changes at different optimization phases, we show in Fig. 6(b) the average number of unique placements executed per episode, where each unique placement refers to a pair of object and receptacle where the agent placed the object on the receptacle. Higher degree of exploration behavior is indicated by higher number of unique placements, and in contrary, lower number of unique placements indicates more exploitation behavior. Fig. 6(b) shows an increase in exploration from LLM-Planner (Base) to LLM-Personalize(IL), partly due to improved plan executability. From IL to ST1, the agent further increased exploration, where it explores placements beyond the behaviors learned from demonstrations. For example, on Scene 3 the average unique placements increased from 2.32 to 9.48. Compared with ST1, ST2 typically shows more exploitation where the agent learns to commit to the correct placement combinations for better task success.
@@ 4.3 Ablation Studies
Table 2: Ablation Study: Cross-domain generalisation success rate. (Boldface: best across the task set, ST1/2: self-training iteration 1 or 2, IL: imitation learning)
@@
5 CONCLUSIONS
@@
We proposed LLM-Personalize, an LLM-based household robotics agent framework capable of long-horizon planning in multi-room, partially observable household scenarios, and an optimization pipeline that personalizes the LLM planner to user preferences. This approach effectively addresses the gap in personalizing LLM planners for household robotics. Our model achieves superior alignment with user preferences, outperforming existing work in the challenging Housekeep rearrangement tasks. Moving forward, the versatility and scalability of the agent design and optimization pipeline makes it a promising solution for a broad range of personalized LLM-powered robotics applications.
6 LIMITATIONS
There are two main limitations in this work. Firstly, although our framework demonstrates significant improvements in aligning LLM planners with human preferences, the current scale of our experiments was constrained by computational resources, (e.g., the LLM API budget). More extensive evaluations involving larger, more diverse household environments could offer further opportunities for enhancing scalability and robustness.
Secondly, while the simulated environment provides valuable insights into the effectiveness of LLM-Personalize, the absence of real-world testing on physical robots and real-world household scenarios leaves room for further validation in dynamic real-world settings. An interesting future direction can aim to address this by deploying our approach on actual robots in household scenarios.
Appendix A Appendix
A.1 Preliminaries: LLMs and Self-training
An LLM Brown et al. (2020) is a transformer-based Vaswani et al. (2017) language model that is capable of performing general-purpose language generation, e.g., GPT Brown et al. (2020), PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023). In practice, users interact with LLMs by sending prompts and receiving responses.
Reinforced Self-Training Iterative Gulcehre et al. (2023) reinforced Self-Training is a sample efficient algorithm for aligning LLMs with human preferences, particularly for the domain of machine translation. The key mechanism involves a grow step where a training dataset is collected by prompting the LLM to generate multiple responses for each prompt, and an improve step, where the dataset is annotated and filtered according to human preferences, followed by fine-tuning the LLM on the dataset using offline RL or supervised learning.
A.2 Controller
Given our primary focus on personalizing LLM planners, we assume a room-level topological map is available (e.g., via semantic mapping Pronobis and Jensfelt (2011)). We make use of the off-the-shelf controller accessible from the Housekeep simulator, which maps each high-level action to a sequence of low-level actions. During a task, the robot navigates and continually updates a room-level topological map using egocentric observations, camera projection matrix, RGBD-aligned pixel-wise instance and semantic masks and relationship sensor to localize objects and receptacles to update the map. The high-level actions from the LLM planner are carried out as follows: 1) for go to, the controller uses the allocentric map and the target entity (obj, rec, room), and executes a sequence of navigation actions to reach the target; 2) for look at, the agent orients itself to face the desired target via look up/down and turn left/right actions; 3) to carry out pick/place, the agent invokes a discrete grab/release action that casts a ray, and if it intersects an obj or rec within 1.5m, it picks or places an object. More details of the controller can be found in Kant et al. (2022).
A.3 Experiment Setup: Baselines
To adapt SayCan to Housekeep with large number of available actions, we adopt the implementation in Song et al. (2023), where the list of affordable high-level actions (e.g., go to a discovered object) are provided in each prompt, together with the list of previously executed actions. For SayPlan, we additionally provide the state and affordance of the receptacles and objects (e.g., pick up) as in Rana et al. (2023) and allow 10 LLM semantic search steps and 5 revision (re-plan) retries for each plan iteration where the final revised plan is executed, and the revision feedback is provided by a verification module (e.g., cannot pick up obj as the agent is holding another object).